OpenStudy (anonymous):

The line 3x-4y=12 is tangent to the parabola y^2=4px find p

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

3x-4y=12 ---> -4y=-3x+12 ---> y = (3/4)x - 3

OpenStudy (solomonzelman):

So the slope of your parabola at the point of tangency is 3/4.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

y^2 = 4px y = ±\(\sqrt{4px}\) y = ±2\(\sqrt{px}\) y\('\) = ±p\(/\sqrt{px}\) Where is the slope 3/4? 3/4 = ±p\(/\sqrt{px}\) and it is above the x axis because the slope is positive, so in ± we take the +. 3/4 = p\(/\sqrt{px}\) 3/(4p) = \(1/\sqrt{px}\) (4p)/3 = \(\sqrt{px}\) 4p = 3√(px) -----> first equation

OpenStudy (solomonzelman):

4p = 3√(px) 4p = 3•√p•√x 4√p = 3√x 16p = 9x 16p / 9 = x ---> point of tangency

OpenStudy (solomonzelman):

oh sorry we choose the below the x-axis part. SO http://www.wolframalpha.com/input/?i=%283%2F4%29%2816p+%2F+9%29+-+3+%3D+-2%E2%88%9A%28p%2816p+%2F+9%29%29

OpenStudy (anonymous):

so what do we do with the point of tangency

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/bj3xaweyae

OpenStudy (solomonzelman):

I solved for p via wolfram and there we see it is once tangent to the p=-9/4 version parabola.